OpenStudy (okdutchman7):

The position function x(t) of a particle moving along an x axis is x=4.0 - 6.0t^2, with x in metres and t in seconds. (a) At what time and (b) where does the particle (momentarily) stop? At what (c) negative time and (d) positive time does the particle pass through the origin? (e) Graph x versus t for the range -5 s to +5 s. (f) To shift the curve rightward on the graph, should we include the term +20t or the term -20t in x(t)? (g) Does that inclusion increase or decrease the value of x at which the particle momentarily stops? I already did (a) and (b) and (e). I need help with the rest.

OpenStudy (osprey):

to me the origin is 0,0. According to the xlsx enclosed the thing doesn't go through the origin ???

OpenStudy (osprey):

modified spreadsheet encl. SHOULD now be three sheets ... plus 20 t. increases

OpenStudy (irishboy123):

(a) At what time .... \( x=4 - 6t^2, \dot x = - 12t = 0\), and \(\dot x = 0 \implies t = 0\), and (b) where \(t = 0 \implies x = 4\) does the particle (momentarily) stop? At what (c) negative time ... \(x = 4 - 6t^2\) so \(x = 0 \implies t = \pm \sqrt{ \dfrac{2}{3} }\) So: \( t = - \sqrt{ \dfrac{2}{3} }\) and (d) positive time does the particle pass through the origin? So: \( t = \sqrt{ \dfrac{2}{3} }\) (e) Graph x versus t for the range -5 s to +5 s.  (f) To shift the curve rightward on the graph, should we include the term +20t or the term -20t in x(t)? to shift rightward write it as: \( x=4 - 6(t-20)^2\) (g) Does that inclusion increase or decrease the value of x at which the particle momentarily stops? Shifting right means particle now stops at \(t = 20\) Check it: \( x=4 - 6(t-20)^2\) \( \dot x=- 2*6(t-20) = 0 \implies t = 20, ~ x = 4 \)

OpenStudy (osprey):

@IrishBoy123 Very elegant expo. and expl. At least I got the graph right !

OpenStudy (irishboy123):

thank you mate !! you can report me for feeding answers, if you like :)

OpenStudy (s4sensitiveandshy):

Done.

OpenStudy (irishboy123):

arsehøle! i still haven't got an email. I might just have to do it myself.

OpenStudy (irishboy123):

https://www.youtube.com/watch?v=GOkZs2RZMQY&list=PLD29CE3C0885891C7&index=10 FWIW: i recommend this, highly. an uber-genius at work. @osprey: Legendre :)

OpenStudy (s4sensitiveandshy):

Offensive..

OpenStudy (irishboy123):

@S4Sensitiveandshy FYI, i found the thing below deeply "offensive". I reported it twice, in each case well before it was announced that this ship had officially hit the iceberg.  Yet still Nada.

OpenStudy (s4sensitiveandshy):

I need glasses to read that

OpenStudy (s4sensitiveandshy):

Gone.

OpenStudy (osprey):

@S4Sensitiveandshy well, yes and no you need glasses. If you save the pic and then read it it's larger. @IrishBoy123 But, then, I think I've seen the post before, so it's there in "larger print" somewhere else.

OpenStudy (irishboy123):

it's legible even on a smartphone. i've tested it and here is a png link https://i.gyazo.com/7e2558c11de97c9727ec5ddf62fb433a.png the original thread is still out there too but the lag here makes it a pain to find.

OpenStudy (irishboy123):

found it: http://openstudy.com/updates/5810b82ae4b075d6916f6fd1

OpenStudy (osprey):

@IrishBoy123 the original thread is still out there too but the lag here makes it a pain to find. Yup; agreed, totally agreed.

OpenStudy (s4sensitiveandshy):

this is my smartphone. http://prnt.sc/df8ny5

OpenStudy (s4sensitiveandshy):

Thanks for the link. I was looking for the thread too, but..i gave up.... because of the lag....

OpenStudy (irishboy123):

thanks @S4Sensitiveandshy .